Short Report – QA/QC Meeting October 2018

Meeting Date: 2018-10-18

Participants / Excused

MG, SB, IF, RM, EPL, TB, IB, PL (8)  / LH, WE, PM

Talking Points


If you are (Tech-)SRP, please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.

SRP & Tech-SRP: please prepare short statement about your site and send to to LH (max. 2 sentences)

Meeting Notes


  • TB showed everyone how to create a fieldbook entry when there is a problem with a device or when maintenance is done.
  • TB also showed us how the SRPs can create a new device on the field book if they need to.
  • There is a new function on GIN, “system group” with which one can combine several devices for a single group, for example for the soil profile.
  • RM had issues creating fieldbook entries when the text was long (software bug). TB suggested dividing the entries into several events. For one event then several devices can be
    added (e.g. cleaning, calibration, etc.)
  • SRPs are going to insert device-related issues in GIN and provide feedback to TB so DO and TB can develop the current options. At some point a link will be established between the notes on the weekly check tables and GIN fieldbooks (TBD with LH).


  1. Automatic fluxes for last 14d (schedule is here)
  2. Online weekly QA/QC checks for nearly all sites
  3. Most recent eddy covariance raw data plots are found in the following folder on our group drive (P) (updated daily):


Site Reports


SRP (LH, fluxes): EC data is complete for the last month except a gap around 25-27 Sep (LANNER did not boot up, see GIN).
Tech-SRP (IB): IB has been working on the mast. He is not sure if changing the height affects the measurements. From now on he will make detailed fieldbook entries on GIN for dates that he is working on the mast so we can check how it affects the measurements.


SRP (IF) / Tech-SRP (PL): Overall, there hasn’t been a gap in the raw data, but in terms of quality after eddy pro calculations some data is kicked out (examples marked below) and hence we have gaps, especially between 12 and 7 AM since 3-4 weeks ago. IF thinks this is due to dew formation on the open path IRGA and others found this sensible since fog is very common in Chamau. In addition, IF has observed that these gaps correlate with higher relative humidity. However, TB suggests that since IRGA is heated this should not happen. The heating of the IRGA cannot be adjusted so IF is going to discuss with WE and keep an eye if this issue persists.


SRP (MG, LH): Recent fluxes shown in the plots were calculated from ICOS csv files instead of the ETH binary files. Fluxes look good, but there is still a bug in sonicread that stops data logging every few weeks and the EC system has to be restarted (remotely).
Tech-SRP: –


SRP (LH): EC data are complete for the last 2 months, no gaps.
Tech-SRP: –


Sonic had to get calibrated hence we have gaps in the data.

SRP (RM): Sonic did not work properly for almost a month and was replaced. New sonic was configured on 9.10.2018 and since then data is complete and fluxes look good.
T-SRP (MS): Licor 7500 was calibrated. Works fine


Eugenie is wondering when the RDS problem will be solved?

SRP (EPL): Fluxes look good. RDS processing still does not work -this needs to be solved soon!
Tech-SRP: –


The sonic partially stopped giving data (similar to what we saw in CH-DAV) that seemed like a regular bug in one of the scripts. Also with regard to the TCP connection
problem WE has written a script to solve the data transfer problem to sonic. TB will discuss with WE to use this script for all sites.

SRP (RM): Unfortunately, we have a data gap of 5 days because the sonic did not work properly. Since 5.10.2018 everything is working again and fluxes look good and complete.
Tech-SRP (MS): Licor 7200 was calibrated. Works fine, drift of CO2 is gone. Camera was moved from field in front of the BAFU Container to the Container itself. Works fine. Camera Settings changed from taken a picture every 4h to every hour, now.



You are here: Short Report – QA/QC Meeting October 2018


Last Updated on 15 Feb 2024 22:56